OpenStudy (anonymous):

Help :) Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. how do I Use temperature conversion formulas to show that these two measures are indeed the same temperature.

sam (.sam.):

°C x 9/5 + 32 = °F

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Always hold this derivation ..\[\frac{ T ^{0}F -32}{ 212-32 }=\frac{ T ^{0}C-0 }{ 100-0 }=\frac{ T ^{0}K -273}{ 373-273 }\]